# Optimistic Thompson Sampling for No-Regret Learning in Unknown Games

### 07 Feb 2024 v1

@@ 3.2 Bandit Feedback
In the bandit feedback setting, Alice can only observe . We propose a framework that combines Optimism algorithm for the stochastic bandits and Noregret algorithm for the full information adversarial bandits. First, we construct a sequence of surrogate full information feedback $\tilde{R}_{t}\in\mathbb{R}^{\mathcal{A}}$ in an optimistic sense, which we refer to as the imagined reward. In particular, we can use an optimism algorithm $E$ for stochastic bandit
$\tilde{R}_{t+1}=E(H_{t+1},Z_{t+1})\in\mathbb{R}^{\mathcal{A}}.$ (4)
Then, we apply a noregret update rule $g_{t}$ to update the sampling distribution with $\tilde{R}_{t+1}$
$X_{t+1}=g_{t}(X_{t},\tilde{R}_{t+1}).$ (5)
This procedure is described in Algorithm 1, termed as Optimism-then-NoRegret (OTN).
Algorithm 1 Optimism-then-NoRegret
@@
6:     Construct reward vector $\tilde{R}_{t+1}$ with an optimism algorithm $E$
7:     Noregret update
@@
The essential part in Algorithm 1 is the construction of the imagined reward vector $\tilde{R}_{t+1}$.
To motivate, we provide a key regret decomposition
@@
For any $a\in\mathcal{A}$, the one-step regret can be decomposed as
$\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right]=(I)+(II)+(III)$
where
$(I)$ $=\mathbb{E}\left[\tilde{R}_{t+1}(a)-\tilde{R}_{t+1}(A_{t})\mid\theta\right],$
$(II)$ $=\mathbb{E}\left[f_{\theta}({a,B_{t}})-\tilde{R}_{t+1}(a)\mid\theta\right],$
$(III)$ $=\mathbb{E}\left[\tilde{R}_{t+1}(A_{t})-f_{\theta}(A_{t},B_{t})\mid\theta\right].$
Summation of $(I)$ reduces to adversarial regret $\Re_{full}(T,\operatorname{adv},\tilde{R})$ for a bounded sequence $\tilde{R}_{t+1}$. For the term $(II)$, sufficient optimism guarantees $\sqrt{T}$-type regret. Therefore, the OTN framework can achieve sublinear regret when equipped with reasonable optimism algorithm and noregret algorithm. Next we discuss several algorithms that can be implied by Algorithm 1.
A celebrated estimation method is importance-weighted estimator (IWE) (Lattimore & Szepesvári, 2020), which combined with Hedge becomes the well known Exp3 algorithm (Auer et al., 2002b) (referred as IWE-Hedge in this paper). Additionally, IWE can also be combined with RM, resulting in IWE-RM for the bandit game. Details of IWE-Hedge and IWE-RM are presented in Appendix D, and their adversarial regrets are provided below.
Proposition 3.2 (Regrets of IWE-Hedge (Auer et al., 2002b)).
Consider play bandit game with IWE-Hedge or IWE-RM, then *$\Re^{*}(T,\text{IWE-Hedge})=\mathcal{O}(\sqrt{T\mathcal{A}\log\mathcal{A}}).$
We also extend the regret matching (RM) (Hart & Mas-Colell, 2000) to our setting under the OTN framework, resulting IWE-RM algorithm (detailed proof can be found in Section D).
Theorem 3.3 (Regret of IWE-RM).
Consider play bandit game with IWE-RM, then *$\Re^{*}(T,\text{IWE-RM})=\mathcal{O}(T^{2/3}\mathcal{A}^{2/3}).$
@@
IWE-RM is our by-product result which is not known in the literature.
However, IWE only requires the bandit reward feedback and does not explicitly utilize the information of opponents’ action. An important idea is that, with the information of opponent’s actions and the knowledge on reward structure $f_{\theta}$, we can estimate $f_{\theta}(a,b)$ for all $(a,b)\in\mathcal{A}\times\mathcal{B}$ by $\tilde{f}_{t+1}(a,b)$ at each time $t$, and construct imagined reward by $\tilde{R}_{t+1}(a)=\min(\tilde{f}_{t+1}(a,B_{t}),1)$. By taking Gaussian distribution as the prior, the mean and variance estimation of $f_{\theta}(a,b)$ up to time $t$ can be recursively updated, denoted as $\mu_{t}(a,b)$ and $\sigma_{t}^{2}(a,b)$. Detailed updating formulas for games in Table 4 are presented in Appendix C. The remaining issue is how to balance exploration and exploitation based on mean and variance estimation.
In stochastic bandit, UCB (Auer et al., 2002a) and TS (Thompson, 1933) are two widely used techniques to achieve optimism. Tailoring to the bandit game, the corresponding constructions using UCB and TS are as follows with $\beta_{t}=\sqrt{2\log{\mathcal{A}\sqrt{t}}}$:
$\begin{cases}\textrm{UCB:}&\tilde{f}_{t+1}(a,B_{t})\>|\>H_{t+1}=\mu_{t}(a,B_{t})+\beta_{t}\sigma_{t}(a,B_{t}),\\ &\tilde{R}_{t+1}(a)=\tilde{f}_{t+1}(a,B_{t})\wedge 1,\forall a\in\mathcal{A}.\\ \textrm{TS:}&\tilde{f}_{t+1}(a,B_{t})\>|\>H_{t+1}\sim N(\mu_{t}(a,B_{t}),\sigma_{t}(a,B_{t})),\\ &\tilde{R}_{t+1}(a)=\tilde{f}_{t+1}(a,B_{t})\wedge 1,\forall a\in\mathcal{A}.\end{cases}$
The effectiveness of UCB combined with Hedge has been studied in (Sessa et al., 2019). Within the OTN framework, UCB can also be combined with RM, resulting in a sublinear regret based on our analysis (see details in Section 4). The regrets of UCB-Hedge and UCB-RM are provided below:
Theorem 3.5 (Regrets of UCB-Hedge and UCB-RM).
Consider playing the bandit feedback game with UCB-Hedge or UCB-RM, then *$\Re^{*}(T,\text{UCB-Hedge})=\mathcal{O}(\sqrt{T\log\mathcal{A}}+\sqrt{\gamma_{T}\beta T}),\ \Re^{*}(T,\text{UCB-RM})=\mathcal{O}(\sqrt{T\mathcal{A}}+\sqrt{\gamma_{T}\beta T}).$
@@
Our regret analysis of UCB-Hedge is under Bayesian setting while that in (Sessa et al., 2019) is under frequentist setting. Due to the setting difference, our regret bound has $\sqrt{\gamma_{T}}$ improvement.
3.3 Failure of TS
We demonstrate that TS combined with RM fails in a class of bandit games by providing a counter example.
@@ 3.3 Failure of TS
Consider a class of matrix games with a payoff matrix $\theta$ defined as
@@
where $\Delta\in(0,1)$. At time $t$, Alice plays action $A_{t}\sim P_{X_{t}}$, and Bob is a best response player who selects action $B_{t}\sim P_{Y_{t}}$ according to the best response strategy $Y_{t}=\min_{y}X_{t}^{\top}\theta y$. Alice receives the noiseless reward $R_{t+1,A_{t},B_{t}}=\theta_{A_{t},B_{t}}$.
Figure 2: Failure probability for TS-RM and OTS-RM.
Intuitively, as long as Alice chooses pure strategy, Bob will exploit Alice, resulting in $\Delta$ regret. Specifically, let $\Omega_{t}$ denote the failure event where Alice selects the 2nd row and Bob chooses the 1st column for all time steps $t^{\prime}\leqslant t$. Assuming uniform initialization for Alice, then $\Re(T)\geqslant 2\mathbb{P}(\Omega_{T})\Delta\cdot T.$ If the event $\Omega_{t}$ occurs with a constant probability for all $t\geqslant 1$, then Alice suffers linear regret.
As shown in Figure 2, TS-RM indeed fails in different setups for hyperparameter $(\Delta,\sigma_{n}^{2})$. Detailed mathematical justification can be found in Appendix E.
Proposition 3.7 (Failure of TS-RM).
Suppose Alice initializes with a uniform strategy and utilizes Regret Matching with TS estimator (TS-RM). For any $\Delta\in(0,1)$ and $\sigma_{n}>0$, there exists a universal constant $c(\Delta,\sigma_{n})>0$ such that for all rounds $t\geqslant 1$, $\mathbb{P}(\Omega_{t})\geqslant c(\Delta,\sigma_{n})$.  [^5] If $\Delta=0.1$ and $\sigma_{n}=0.1$, we have $c(\Delta,\sigma_{n})\approx 0.54$.
The proof of Proposition 3.7 can be found in Appendix E.
3.4 A Simple Fix: Optimistic Sampling
By taking the maximum of multiple functions sampled from the posterior, we can improve the probability of being optimistic for the estimator. At time $t$, after observing $H_{t+1}=(H_{t},A_{t},B_{t},R_{t+1,A_{t},B_{t}})$, we can perform optimistic sampling by first sampling $M_{t+1}$ independent normal random samples and constructing the estimator: $\forall a\in\mathcal{A}$
$\tilde{f}^{\operatorname{TS},j}_{t+1}(a,B_{t})$ $\sim N(\mu_{t}(a,B_{t}),\sigma_{t}(a,B_{t})),j\in[M_{t+1}]$
$\tilde{f}^{\operatorname{OTS}}_{t+1}(a,B_{t})$ $:=\max_{j\in[M_{t+1}]}\tilde{f}^{\operatorname{TS},j}_{t+1}(a,B_{t}),$
$\tilde{R}^{\operatorname{OTS}}_{t+1}(a)$ $:=\operatorname{clip}_{[0,1]}\left(\tilde{f}_{t+1}^{\operatorname{OTS}}(a,B_{t})\right).$
To intuitively understand why optimistic samples help, let us recall Example 3.1. Denote $\tilde{R}^{OTS}_{t+1}(1st)$ and $\tilde{R}^{OTS}_{t+1}(2nd)$ as the optimistic reward for the first and second row actions, respectively. Then the probability of $\tilde{R}^{OTS}_{t+1}(1st)>\tilde{R}^{OTS}_{t+1}(2nd)$ increases as $M_{t+1}$ grows, this implies a decay of $\mathbb{P}(\Omega_{t})$ over time $t$. Detailed mathematical justification can be found in Appendix E. Such a simple optimistic version of TS (referred to as OTS) has the following regret guarantee.
Theorem 3.8 (Regrets of OTS-Hedge and OTS-RM).
For any full information adversarial bandit $\operatorname{adv}$ working with the imagined reward sequence $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t+1},t=0,1,\ldots)$ constructed by OTS with $M_{1}=M_{2}=\ldots=M_{T}=M=\mathcal{O}(\log\mathcal{A}T)$, the combined algorithm $\pi=\pi^{\operatorname{adv-OTS}}$ enjoys the regret bound:
*$\Re^{*}(T,\pi,\theta)\leqslant\Re_{\operatorname{full}}(T,\operatorname{adv},\tilde{R}^{\operatorname{est}})+\sqrt{\log(\mathcal{A}T)I(\theta;H_{T})T}$
where $\Re_{\operatorname{full}}$ is defined in Equation 3.
Note that as long as the imagined $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t+1},t=0,1,\ldots)$ satisfies bounded property $\tilde{R}_{t}\in[0,1]^{\mathcal{A}}$, then if we take $\operatorname{adv}$ to be Hedge, $\Re_{\operatorname{full}}(T,\text{Hedge},\tilde{R}^{\operatorname{est}})=\mathcal{O}(\sqrt{T\log\mathcal{A}})$. And if taking $\operatorname{adv}$ to be RM, then $\Re_{\operatorname{full}}(T,\text{RM},\tilde{R}^{\operatorname{est}})=\mathcal{O}(\sqrt{T\mathcal{A}})$.
@@ 3.4 A Simple Fix: Optimistic Sampling
The regret bound for UCB shares the same order as OTS but there is a fundamental difference, i.e., OTS actually is a stochastic upper confidence bound method. Further, from practical implementation view, OTS has some potential benefits like obtaining an (approximate) posterior sample of complex function is computationally easier than compute the UCBs of the function.
@@ 4 Regret Analysis
Our analysis road map is as follows. First, recall that Proposition 3.1 presents a general regret decomposition
$\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right]=(I)+(II)+(III)$
Summation of $(I)$ reduces to adversarial regret $\Re_{full}(T,\operatorname{adv},\tilde{R})$ for a bounded sequence $\tilde{R}_{t+1}$. For the term $(II)$, the following definition of optimism is a sufficient condition for $\sqrt{T}$-type regret. In addition, Proposition 3.1 also applies to IWE and UCB, resulting in our by-product regret bounds of IWE-RM and UCB-RM (see Table 2 and details are presented in Appendix F).
@@
We say the constructed imagined reward sequence $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t},t\in\mathbb{Z}_{++})$ is optimistic if for any action $a\in\mathcal{A}$, $\mathbb{P}(f_{\theta}(a,B_{t})\geqslant\tilde{R}_{t+1}(a)\mid\theta)\leqslant\mathcal{O}(1/\sqrt{T}).$
OTS can satisfy Definition 4.1 by selecting proper $M_{t+1}$ whereas TS cannot satisfy. UCB sequence can also satisfy Definition 4.1. The term $(III)$ can be bounded by one-step information gain $I(\theta;R_{t+1,A_{t},B_{t}}|H_{t})$ using differential entropy of Gaussian distribution. Here, we give a general regret bound for any imagined reward sequence.
@@
For any full information adversarial bandit $\operatorname{adv}$ working the imagined reward $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t},t\in\mathbb{Z}_{++})$ constructed by estimation algorithm $\operatorname{est}$, if the $\tilde{R}^{\operatorname{est}}$ satisfy Definition 4.1, the regret of the combined algorithm $\pi=\pi^{\operatorname{adv-est}}$ enjoys
@@
where $\Re_{\operatorname{full}}$ is defined in Equation 3 and $\beta=\mathcal{O}(\log\mathcal{A}T)$.
Proof sketch. First, we derive a general regret decomposition for any imagined reward sequence, and bound each term separably. We bound the first term via a reduction to adversarial regret. For the second term, we introduce the generic UCB sequence and lower confidence bound (LCB) sequence and derive a general regret bound for these sequences. We bound the third term via information-theoretic quantity. Finally, combining the three bounds together yields the regret upper bound. The detailed proof is provided in Appendix F.
To characterize how much uncertainty reduction is in a particular game structure when observing new information, we define the maximum information gain.
@@
$\gamma_{T}:=\max_{A_{0:T},B_{0:T}}I(\theta;A_{0},B_{0},\ldots,A_{T-1},B_{T-1})$
where $I(X;Y)$ is the mutual information between random variables $X$ and $Y$.
@@
An important property of the Gaussian distribution is that the information gain does not depend on the observed rewards. This is because the posterior covariance of a multivariate Gaussian is a deterministic function of the sampled points. For this reason, this maximum information ratio $\gamma_{T}$ in Definition 4.3 is well defined. That is, $I(\theta;H_{T})=I(\theta;A_{0},B_{0},\ldots,A_{T-1},B_{T-1})\leqslant\gamma_{T}.$
We adopt the results from (Srinivas, Krause, Kakade, and Seeger, 2009) which gives the bounds of $\gamma_{T}$ for a range of commonly used covariance functions: finite dimensional linear, squared exponential and Matern kernels, whose details can be found in Appendix F.
@@
Utilizing the information of opponent’s actions and the underlying reward structure can help resolve the curse of multi-agent. For example, if the reward structure can be model by squared exponential kernels, the final regret of OTS-Hedge is $\mathcal{O}\left(\left(\sqrt{\log\mathcal{A}}+\sqrt{\log(\mathcal{A}T)\log(T)^{d+1}}\right)\sqrt{T}\right)$. In comparison to the standard adversarial multi-armed bandit regret bound $\mathcal{O}(\sqrt{T\mathcal{A}\log\mathcal{A}})$ (e.g., EXP3 do not have the information of opponent’s action), this regret bound does NOT depend polynomially on the number of actions $\mathcal{A}$ and the number of opponent’s action $\mathcal{B}$, similarly to the ideal full-information setting. This is possible due to the information of opponent action which helps capturing the underlying structure of the reward function that depends on actions of all agents.

### 25 Feb 2024 v2

@@ 3.2 Bandit Feedback
In the bandit feedback setting, Alice can only observe a noisy version of the reward for the action she selects. We propose a framework that combines an optimism algorithm for stochastic bandits with a no-regret algorithm for full information adversarial bandits. First, we construct a sequence of surrogate full information feedback $\tilde{R}_{t}\in\mathbb{R}^{\mathcal{A}}$ in an optimistic sense, which we refer to as the imagined reward. Specifically, we use an optimistic estimation algorithm $E$ to construct reward vector $\tilde{R}_{t+1}=E(H_{t+1},Z_{t+1})\in\mathbb{R}^{\mathcal{A}}.$ Then, we apply a no-regret update rule $g_{t}$ to update the sampling distribution with $\tilde{R}_{t+1}$ as $X_{t+1}=g_{t}(X_{t},\tilde{R}_{t+1}).$ This procedure is described in Algorithm 1, termed Optimism-then-NoRegret (OTN).
Algorithm 1 Optimism-then-NoRegret (OTN)
@@
6:     Construct the reward vector $\tilde{R}_{t+1}$ with an optimistic estimation algorithm $E$
7:     No-regret update
@@
The essential part of Algorithm 1 is the construction of the imagined reward vector $\tilde{R}_{t+1}$.
To elucidate the strategic underpinnings of adversarial bandit games, we introduce a comprehensive regret decomposition. This analytical framework sheds light on the subtleties of strategic decision-making against adversaries.
@@
Given any action $a\in\mathcal{A}$, we can dissect the one-step regret as follows:
$\mathbb{E}\left[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}|\theta\right]=(I)+(II)+(III),$
where:
$(I)$ $=\mathbb{E}\left[\tilde{R}_{t+1}(a)-\tilde{R}_{t+1}(A_{t})|\theta\right],$
$(II)$ $=\mathbb{E}\left[f_{\theta}(a,B_{t})-\tilde{R}_{t+1}(a)|\theta\right],$
$(III)$ $=\mathbb{E}\left[\tilde{R}_{t+1}(A_{t})-f_{\theta}(A_{t},B_{t})|\theta\right].$
The aggregation of term $(I)$ essentially quantifies the adversarial regret $\mathcal{R}_{\text{full}}(T,\operatorname{adv},\tilde{R})$ within a bounded sequence $\tilde{R}_{t+1}$. Through the application of sufficient optimism, term $(II)$ and $(III)$ enables the realization of $\sqrt{T}$-type regret minimization. Thus, the OTN framework facilitates the attainment of sublinear regret, contingent upon the integration of well-designed optimism and no-regret algorithms. We now proceed to examine various algorithms that seamlessly integrate within the Algorithm 1 framework.
The importance-weighted estimator (IWE) (Lattimore & Szepesvári, 2020) stands as a cornerstone in the realm of bandit algorithms. When amalgamated with Hedge, it forms the basis of the renowned Exp3 algorithm (Auer et al., 2002b; Kocák et al., 2014; Neu, 2015), hereafter referred to as IWE-Hedge. Furthermore, the integration of IWE with Regret Matching (RM) introduces the IWE-RM strategy for bandit games. The nuances of both IWE-Hedge and IWE-RM, alongside their implications for adversarial regret, are meticulously outlined in Appendix D.
Proposition 3.2 (IWE-Hedge (Exp3) Analysis).
Engaging in a bandit game utilizing the IWE-Hedge strategy results in *$\Re^{*}(T,\text{IWE-Hedge})=\mathcal{O}(\sqrt{T|\mathcal{A}|\log|\mathcal{A}|}).$
The adaptation of Regret Matching (RM) (Hart & Mas-Colell, 2000) through the lens of IWE under bandit feedback culminates in the innovative IWE-RM algorithm.
Theorem 3.3 (Analysis of IWE-RM).
Implementing the IWE-RM approach in bandit gameplay yields *$\Re^{*}(T,\text{IWE-RM})=\mathcal{O}(T^{2/3}|\mathcal{A}|^{2/3}).$
@@
The derivation of IWE-RM represents a novel contribution, previously unexplored in the literature. Detailed justification is available in Section D.
Crucially, while IWE hinges on bandit reward feedback, it does not inherently leverage opponent action information. By incorporating opponent action insights and reward structure knowledge $f_{\theta}$, we can refine our estimation strategies. Specifically, $f_{\theta}(a,b)$ for all pairs $(a,b)\in\mathcal{A}\times\mathcal{B}$ is approximated by $\tilde{f}_{t+1}(a,b)$, facilitating the construction of an imagined reward $\tilde{R}_{t+1}(a)$. Adopting a Gaussian distribution as the prior enhances the precision of mean $\mu_{t}(a,b)$ and variance $\sigma_{t}^{2}(a,b)$ estimations over time. The articulation of these estimation processes, alongside strategies for balancing exploration and exploitation, is detailed in Appendix C.
In the stochastic bandit landscape, the Upper Confidence Bound (UCB) (Auer et al., 2002a) and Thompson Sampling (TS) (Thompson, 1933) emerge as pivotal methods for instilling optimism in algorithmic choices. These methods are adeptly tailored to bandit games, with specific constructions for UCB and TS elucidated below, employing a parameter $\beta_{t}=\sqrt{2\log{|\mathcal{A}|\sqrt{t}}}$:
$\begin{cases}\text{UCB:}&\tilde{f}_{t+1}(a,B_{t})|H_{t+1}=\mu_{t}(a,B_{t})+\beta_{t}\sigma_{t}(a,B_{t}),\\ &\tilde{R}_{t+1}(a)=\min(\tilde{f}_{t+1}(a,B_{t}),1),\forall a\in\mathcal{A}.\\ \text{TS:}&\tilde{f}_{t+1}(a,B_{t})|H_{t+1}\sim\mathcal{N}(\mu_{t}(a,B_{t}),\sigma_{t}^{2}(a,B_{t})),\\ &\tilde{R}_{t+1}(a)=\min(\tilde{f}_{t+1}(a,B_{t}),1),\forall a\in\mathcal{A}.\end{cases}$
The synergy of UCB with Hedge, and its integration within the OTN framework, underscores a pathway to sublinear regret, as substantiated through our analytical endeavors (details in Section 4). The comparative analysis of UCB-Hedge and UCB-RM, delineated below, highlights the efficacy of these approaches:
Theorem 3.5 (Efficacy of UCB-Hedge and UCB-RM).
Application of UCB-Hedge or UCB-RM in the context of bandit feedback games ensures *$\Re^{*}(T,\text{UCB-Hedge})=\mathcal{O}(\sqrt{T\log|\mathcal{A}|}+\sqrt{\gamma_{T}\beta T}),$ and similarly, *$\Re^{*}(T,\text{UCB-RM})=\mathcal{O}(\sqrt{T|\mathcal{A}|}+\sqrt{\gamma_{T}\beta T}).$
@@
The analytical framework for UCB-Hedge adopts a Bayesian perspective, contrasting with the frequentist approach taken in existing studies (Sessa et al., 2019). This methodological divergence yields a $\sqrt{\gamma_{T}}$ improvement in our regret bounds.
3.3 Challenges with TS
We explore the inefficacy of integrating Thompson Sampling (TS) with Regret Matching (RM) in certain bandit game contexts through a demonstrative counterexample.
@@ 3.3 Challenges with TS
Consider matrix games characterized by a payoff matrix $\theta$:
@@
with $\Delta\in(0,1)$. In each round $t$, Alice selects her action $A_{t}\sim P_{X_{t}}$, while Bob, employing a best response strategy, chooses $B_{t}\sim P_{Y_{t}}$, aiming to maximize his payoff against Alice’s choice. The reward for Alice, $R_{t+1,A_{t},B_{t}}=\theta_{A_{t},B_{t}}$, is determined without noise.
Figure 2: Illustration of failure probabilities for TS-RM versus OTS-RM strategies.
This setup reveals Bob’s ability to exploit Alice if she adheres to a static strategy, leading to a regret increment proportional to $\Delta$. Define $\Omega_{t}$ as the scenario where Alice consistently chooses the second row, and Bob the first column, across all iterations. Given Alice’s strategy begins uniformly, $\Re(T)\geqslant 2\mathbb{P}(\Omega_{T})\Delta\cdot T$, suggesting potential for linear regret if $\Omega_{t}$ persists with consistent probability.
Proposition 3.7 (Limitations of TS-RM).
When Alice employs a uniform strategy and integrates TS with RM, for any $\Delta\in(0,1)$ and noise variance $\sigma_{n}>0$, there exists a constant $c(\Delta,\sigma_{n})>0$ ensuring $\mathbb{P}(\Omega_{t})\geqslant c(\Delta,\sigma_{n})$ for all $t\geqslant 1$, highlighting sustained linear regret.
3.4 Mitigating Strategy: Optimistic Sampling
Optimistic sampling, leveraging multiple posterior samples, emerges as a strategic remedy to enhance the estimator’s optimism. Upon observing $H_{t+1}=(H_{t},A_{t},B_{t},R_{t+1,A_{t},B_{t}})$, optimistic sampling involves generating $M_{t+1}$ independent normal samples to construct an optimistic estimator for each action $a\in\mathcal{A}$:
$\tilde{f}^{\operatorname{TS},j}_{t+1}(a,B_{t})$ $\sim\mathcal{N}(\mu_{t}(a,B_{t}),\sigma_{t}(a,B_{t})),\forall j\in[M_{t+1}],$
$\tilde{f}^{\operatorname{OTS}}_{t+1}(a,B_{t})$ $=\max_{j\in[M_{t+1}]}\tilde{f}^{\operatorname{TS},j}_{t+1}(a,B_{t}),$
$\tilde{R}^{\operatorname{OTS}}_{t+1}(a)$ $=\text{clip}_{[0,1]}\left(\tilde{f}_{t+1}^{\operatorname{OTS}}(a,B_{t})\right).$
Reflecting on Example 3.1, implementing OTS alters the probability dynamics favorably over time, suggesting a reduction in the likelihood of $\Omega_{t}$. This adjustment confirms OTS’s efficacy in countering the limitations of TS in adversarial settings.
Theorem 3.8 (Advantages of OTS).
Incorporating OTS alongside any full-information adversarial bandit strategy $\operatorname{adv}$ yields an enhanced regret bound:
*$\Re^{*}(T,\pi,\theta)\leqslant\Re_{\operatorname{full}}(T,\operatorname{adv},\tilde{R}^{\operatorname{OTS}})+\sqrt{\log(|\mathcal{A}|T)I(\theta;H_{T})T},$
underscoring the strategic benefit of integrating optimism into the estimation process.
@@ 3.4 Mitigating Strategy: Optimistic Sampling
The efficacy of the OTS-type method hinges on the bounded nature of the imagined rewards, $\tilde{R}^{\operatorname{OTS}}=(\tilde{R}^{\operatorname{OTS}}_{t+1},t=0,1,\ldots)$, where each $\tilde{R}_{t}$ lies within the range $[0,1]^{\mathcal{A}}$. This constraint ensures that the reward estimations do not exceed plausible limits, thereby maintaining the integrity of the learning process. When the adversarial bandit algorithm $\operatorname{adv}$ is instantiated as Hedge, the regret bound under the OTS framework, $\Re_{\operatorname{full}}(T,\text{Hedge},\tilde{R}^{\operatorname{OTS}})$, scales optimally as $\mathcal{O}(\sqrt{T\log|\mathcal{A}|})$, demonstrating efficiency in a wide array of strategic scenarios. Similarly, if $\operatorname{adv}$ is implemented as RM, the regret bound, $\Re_{\operatorname{full}}(T,\text{RM},\tilde{R}^{\operatorname{OTS}})$, achieves a comparable rate of $\mathcal{O}(\sqrt{T|\mathcal{A}|})$, signifying robustness across varying game dynamics. This adaptability is a testament to the versatility and practical utility of the OTS strategy in navigating the complexities of unknown game environments.
Remark 3.10.
While UCB and OTS both aim to augment algorithmic performance through optimism, OTS distinguishes itself by adopting stochastic bounds, offering practical advantages in complex environments where sampling efficiency and adaptability surpass traditional UCB computations.
@@ 4 Regret analysis
Our analysis roadmap is outlined as follows. First, we recall Proposition 3.1, which provides a general framework for regret decomposition:
$\mathbb{E}[R_{t+1,a,B_{t}}-R_{t+1,A_{t},B_{t}}|\theta]=(I)+(II)+(III).$
The summation of term $(I)$ leads to the reduction of adversarial regret $\Re_{\text{full}}(T,\operatorname{adv},\tilde{R})$ for a sequence of bounded imagined rewards $\tilde{R}_{t+1}$. For term $(II)$, we define sufficient optimism as a condition to ensure $\sqrt{T}$-type regret. Furthermore, Proposition 3.1 extends to both IWE and UCB, resulting in the derivation of by-product regret bounds for IWE-RM and UCB-RM, as detailed in Table 1 and further elaborated in Appendix F.
@@
The constructed imagined reward sequence $\tilde{R}^{\text{est}}=(\tilde{R}_{t},t\in\mathbb{Z}_{++})$ is deemed optimistic if for any action $a\in\mathcal{A}$, it holds that
$\mathbb{P}(f_{\theta}(a,B_{t})\geqslant\tilde{R}_{t+1}(a)|\theta)\leqslant\mathcal{O}(1/\sqrt{T}).$
OTS adheres to Definition 4.1 by appropriately selecting $M_{t+1}$, whereas TS does not. The UCB sequence also satisfies this optimism criterion. Term $(III)$ can be bounded by the one-step information gain $I(\theta;R_{t+1,A_{t},B_{t}}|H_{t})$, utilizing the differential entropy of Gaussian distributions.
@@
For any full information adversarial bandit $\operatorname{adv}$, utilizing an imagined reward sequence $\tilde{R}^{\operatorname{est}}=(\tilde{R}_{t},t\in\mathbb{Z}_{++})$ constructed by an estimation algorithm $\operatorname{est}$ that satisfies Definition 4.1, the combined algorithm $\pi=\pi^{\operatorname{adv-est}}$ enjoys the regret bound
@@
where $\Re_{\operatorname{full}}$ is as defined in Definition 2, and $\beta=\mathcal{O}(\log|\mathcal{A}|T)$.
Proof sketch. We start with a general regret decomposition for any imagined reward sequence, bounding each term separately. The first term is bounded by reducing it to adversarial regret. For the second term, we utilize generic UCB and LCB sequences to derive a general regret bound. The third term is bounded using an information-theoretic quantity. Combining these bounds yields the overall regret upper bound, with detailed proofs available in Appendix F.
To quantify the uncertainty reduction in specific game structures upon observing new information, we introduce the maximum information gain:
@@
$\gamma_{T}:=\max_{A_{0:T},B_{0:T}}I(\theta;A_{0},B_{0},\ldots,A_{T-1},B_{T-1}),$
where $I(X;Y)$ denotes the mutual information between random variables $X$ and $Y$.
@@
A notable property of Gaussian distributions is that the information gain does not depend on the observed rewards, as the posterior covariance of a multivariate Gaussian is determined solely by the sampled points. Consequently, the maximum information ratio $\gamma_{T}$ in Definition 4.3 is well-defined, implying that $I(\theta;H_{T})=I(\theta;A_{0},B_{0},\ldots,A_{T-1},B_{T-1})\leqslant\gamma_{T}$.
The bounds on $\gamma_{T}$ for various commonly used covariance functions, including finite-dimensional linear, squared exponential, and Matern kernels, are derived from (Srinivas et al., 2009) and detailed in Appendix F.
@@
The integration of information on opponent action and reward structures emerges as a potent strategy to counteract the challenges inherent in multi-agent settings, often referred to as the curse of multi-player. Notably, when employing squared exponential kernels to model the reward dynamics, the regret for OTS-Hedge is refined to $\mathcal{O}\left(\left(\sqrt{\log|\mathcal{A}|}+\sqrt{\log(|\mathcal{A}|T)\log(T)^{d+1}}\right)\sqrt{T}\right)$, a stark contrast to the polynomial dependence on the action spaces $\mathcal{A}\times\mathcal{B}$ observed in traditional approaches. This deviation from conventional adversarial algorithms, exemplified by the EXP3’s regret bound of $\mathcal{O}(\sqrt{T\mathcal{A}\log\mathcal{A}})$, underscores the inefficiency of models that neglect opponent actions. By leveraging such information, OTS-Hedge not only alleviates the curse of multi-player but also significantly surpasses the limitations of standard adversarial bandit methods, echoing the benefits of a full-information framework. This paradigm shift, driven by the strategic exploitation of opponent actions and the nuanced understanding of reward structures, marks a significant leap towards deciphering the complexities of multi-player environments and setting a new benchmark for adversarial strategies.